OpenStudy (will.h):

Good morning yall quick question Alg 2

OpenStudy (will.h):

The last 2 options are \[4\cos(x - \frac{ \pi }{ 2 })\] \[-4\cos(x - \frac{ \pi }{ 2 })\]

OpenStudy (will.h):

@IrishBoy123 Can you help

OpenStudy (irishboy123):

if you want to figure it out yourself, just set x = 0 and \(x = \pi /2\) and see which fits it has to be one of he last two because setting x = 0 clearly rules out the sin ones

OpenStudy (will.h):

it is C correct? @IrishBoy123 i need this right please :)

OpenStudy (irishboy123):

have a look at d \(-4\cos(\pi/2 - \pi/2) = - 4 \cos 0 = -4\) make sure you are happy with answer!

OpenStudy (will.h):

Thank you <3 i have 1 more if it is fine ;)